# On priors and scale cuts in EFT-based full-shape analyses

### v1

@@ 4.1 PT Challenge analysis
Figure 2: Posterior distributions from the different analyses of the PT challenge simulation for the WC model (upper left panel), the EC1 model (upper right panel), the EC2 model (lower left panel) and the EC3 model (lower right panel). The results are inferred from the four BOSS-like data chunks (green), as well as from the single snapshot at $z_{\rm eff}=0.61$ with the covariance matching the BOSS survey volume (blue), and the covariance rescaled to match the full PT challenge volume 566 ($h^{-1}$pc)3 (red). For the EC2 model, we also show the analysis with $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ and the covariance corresponding to the full simulation volume (in pink).
@@
All in all, we conclude that the WC model is somewhat restrictive. First, the posterior of $\omega_{cdm}$ is affected by the priors generated by the inaccuracy of the WC stochasticity model. Second, the WC model applies overoptimistic scale cuts in the $z_{3}$ redshift bin which significantly bias parameter recovery, overestimating $\sigma_{8}$ by more than $5\%$. Unlike the Bayesian marginalization effect, this is a genuine systematic bias that does not vanish with better priors or larger data volumes. In contrast, the EC3 model leads to a small ($<1.5\%$) systematic bias in the recovery of cosmological parameters. This bias constitutes a small negligible fraction of the statistical error in the actual BOSS data. Using more observables, like the cross-correlation with CMB lensing or imposing informative simulation-based priors on nuisance parameters from Ivanov et al. (2024b) can mitigate the Bayesian marginalization effects present in the EC3 analysis with the BOSS volume.
@@ 4.2 BOSS-like analysis
$\Delta\omega_{cdm}/\omega_{cdm}$ $-0.034_{-0.047}^{+0.041}$ $-0.018_{-0.005}^{+0.005}$ $0.004_{-0.060}^{+0.052}$ $-0.0098_{-0.0065}^{+0.0061}$ $-0.006_{-0.054}^{+0.048}$ $0.0088_{-0.0047}^{+0.0043}$
$\Delta h/h$ $-0.001_{-0.018}^{+0.017}$ $-0.0028_{-0.0014}^{+0.0014}$ $0.004_{-0.020}^{+0.019}$ $-0.0015_{-0.0017}^{+0.0016}$ $0.003_{-0.019}^{+0.017}$ $0.0014_{-0.0014}^{+0.0014}$
$\Delta\sigma_{8}/\sigma_{8}$ $-0.031_{-0.058}^{+0.054}$ $0.049_{-0.005}^{+0.005}$ $-0.060_{-0.061}^{+0.053}$ $0.046_{-0.005}^{+0.005}$ $-0.082_{-0.047}^{+0.041}$ $0.015_{-0.005}^{+0.005}$
@@
Let us comment on the large bias that we have seen in the WC model. A natural question is why this has not been identified before in the analyses of D’Amico et al. (2019, 2024)? These studies mostly focused on the Nseries simulations whose volume is smaller than PT Challenge and which implements the effects of the survey selection function and survey geometry. The Nseries mocks were also analyzed in detail with the pipeline associated with the CLASS-PT code in  Chudaykin et al. (2021a); Philcox and Ivanov (2022); Ivanov et al. (2023). These papers found an excellent agreement between analyses based on PT Challenge and Nseries, which served as a validation of the window-free method of Philcox (2021a, b) used in these works. The only difference w.r.t. the studies based on the PyBird code is that they used the standard forward model for the survey geometry, as opposed to the window-free method. This suggests that the apparent tension between the Nseries analyses of WC and EC might stem from inaccuracies in the window function treatment used by the WC team.
@@
Acknowledgments. MI thanks George Efstathiou for inspiring discussions. We thank Guido D’Amico for encouraging us to publish this note. All numerical calculations have been performed with the Helios cluster at the Institute for Advanced Study, Princeton. This work was supported in part by MEXT/JSPS KAKENHI Grant Numbers JP20H05861, JP21H01081, JP22K03634, JP24H00215 and JP24H00221.

### v2

@@ 4.1 PT Challenge analysis
Figure 2: Posterior distributions from the different analyses of the PT challenge simulation for the WC model (upper left panel), the EC1 model (upper right panel), the EC2 model (lower left panel) and the EC3 model (lower right panel). The results are inferred from the four BOSS-like data chunks (green), as well as from the single snapshot at $z_{\rm eff}=0.61$ with the covariance matching the BOSS survey volume (blue), and the covariance rescaled to match the full PT challenge volume 566 ($h^{-1}$Gpc)3 (red). For the EC2 model, we also show the analysis with $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ and the covariance corresponding to the full simulation volume (in pink).
@@
All in all, we conclude that the WC model is somewhat restrictive. First, the posterior of $\omega_{cdm}$ is affected by the priors generated by the inaccuracy of the WC stochasticity model. Second, the WC model applies overoptimistic scale cuts in the $z_{3}$ redshift bin which significantly bias parameter recovery, overestimating $\sigma_{8}$ by more than $5\%$. Unlike the Bayesian marginalization effect, this is a genuine systematic bias that does not vanish with better priors or larger data volumes. In contrast, the EC3 model leads to a small ($1.5\%$) systematic bias in the recovery of cosmological parameters. This bias constitutes a small negligible fraction of the statistical error in the actual BOSS data. Using more observables, like the cross-correlation with CMB lensing or imposing informative simulation-based priors on nuisance parameters from Ivanov et al. (2024b) can mitigate the Bayesian marginalization effects present in the EC3 analysis with the BOSS volume.
@@ 4.2 BOSS-like analysis
$\Delta\omega_{cdm}/\omega_{cdm}$ $034_{-0.047}^{+0.041} 0.0$$18_{-0.005}^{+0.005} 0.0$$04_{-0.060}^{+0.052} 0.0$$098_{-0.0065}^{+0.0061} 0.0$$06_{-0.054}^{+0.048} 0.0$$0088_{-0.0047}^{+0.0043}$
$\Delta h/h$ $001_{-0.018}^{+0.017} 0.0$$028_{-0.0014}^{+0.0014} 0.0$$04_{-0.020}^{+0.019} 0.0$$015_{-0.0017}^{+0.0016} 0.0$$03_{-0.019}^{+0.017} 0.0$$0014_{-0.0014}^{+0.0014}$
$\Delta\sigma_{8}/\sigma_{8}$ $031_{-0.058}^{+0.054} 0.0$$49_{-0.005}^{+0.005} 0.0$$60_{-0.061}^{+0.053} 0.0$$46_{-0.005}^{+0.005} 0.0$$82_{-0.047}^{+0.041} 0.0$$015_{-0.005}^{+0.005}$
@@
Let us briefly summarize the previous tests of the WC analysis pipeline performed in the literature. As a first step, the Refs. D’Amico et al. (2019, 2024) validated the analysis pipeline using the Patchy Mocks Kitaura et al. (2016), which did not show any significant bias in parameter recovery. This is expected, as the Patchy Mocks use an approximate gravity solver and thus can not reliably predict clustering on mildly non-linear scales.  [^14] The Patchy Mocks underestimate the power of the power spectrum monopole at $k\sim 0.2\,h\text{Mpc}^{-1}$ compared to that measured from the BOSS DR12 data Kitaura et al. (2016). Secondly, the Refs. D’Amico et al. (2024) analyzed the mean of 84 Nseries mocks Alam et al. (2017b). These mocks are based on full N-body simulations and implement the effects of the survey selection function and survey geometry. They reported a $4.7\%$ bias in the $\sigma_{8}$ recovery and a $1.3\%$ lower bias in the $\omega_{cdm}$ recovery from the power spectrum using their pipeline in complete agreement with our findings. That said, the interpretations of the results are different. Ref. D’Amico et al. (2024) quantifies the parameter shifts in terms of the statistical uncertainty from the analysis of the CMASS NGC data chunk, while here we compare these shifts to the statistical error from the actual BOSS data analysis (including all four BOSS data chunks). Clearly, the significance of the bias in parameter recovery when expressed in units of the standard deviation is larger in our case than in the analysis of D’Amico et al. (2024).
@@
5 Conclusion
In this work, we explore the different analysis choices associated with the CLASS-PT and PyBird codes. Analysis settings fall into two main categories: EFT modeling (EFT parameters and priors) and data cuts. Below, we outline the implications of these settings for a realistic BOSS-like analysis.
The EFT modeling of the two fitting pipelines differs in several aspects. First, the EC model includes two additional EFT parameters: the next-to-leading order $k^{4}$ counterterm $\tilde{c}$ and the real space scale-dependent stochastic noise counterterm $a_{0}$. Second, the priors on EFT parameters vary significantly between the EC and WC pipelines, with the WC model employing tighter priors than those used in the EC model. However, we find that these differences have only a marginal impact on the resulting parameter constraints, which validates the prior choices made in both analyses. A final and more important difference is that the WC model neglects the scale-dependent stochastic contribution to the power spectrum monopole, which leads to a $0.8\sigma$ downward shift in $\omega_{cdm}$ in the realistic BOSS-like analysis. In contrast, the EC model consistently incorporates scale-dependent stochastic bias and yields an unbiased inference of $\omega_{cdm}$. Thus, it is important to use a consistent model for stochasticity, which requires including a $k^{2}$ stochastic counterterm in both the monopole and quadrupole.
Another important aspect of EFT-based full-shape analyses is the choice of the momentum-scale cut $k_{\rm max}$. Using overoptimistic scale cuts amplifies theory-driven systematic error, which can significantly bias cosmological constraints. The WC model adopts $k_{\rm max}=0.25\,h\text{Mpc}^{-1}$ in the $z_{3}$ redshift bin, which significantly biases parameter recovery, overestimating $\sigma_{8}$ by $\approx 5\%$ (equivalent to $1\sigma$), or by $\approx 4\%$ when using the $k_{\rm max}=0.23\,h\text{Mpc}^{-1}$ from Refs. D’Amico et al. (2019); Colas et al. (2020). By contrast, the EC team employs a more conservative data cut, $k_{\rm max}=0.20\,h\text{Mpc}^{-1}$, for the same data chunk, which results in only a small ($1.5\%$) systematic bias in the recovery of cosmological parameters. This occurs because the theoretical error associated with the two-loop corrections grows rapidly with $k_{\rm max}$, so that even a modest increase of $k_{\rm max}$ from $0.2~\,h\text{Mpc}^{-1}$ to $0.25~\,h\text{Mpc}^{-1}$ (or $0.23~\,h\text{Mpc}^{-1}$) significantly amplifies the error, making it a significant source of systematic uncertainty.  [^16] Using a larger value of $k_{\rm max}$ in the WC analysis alleviates prior volume projection effects, but at the cost of increasing the theory-driven systematic bias in parameter recovery. We conclude that the difference in scale cuts is the main source of the mismatch between the EC and WC analyses, a point that has been largely overlooked in the literature. This discrepancy in $k_{\rm max}$ choices also produces differences in non-Gaussianity constraints between the EC Cabass et al. (2022a, b) and WC D’Amico et al. (2022) analyses, which has been later confirmed in studies of the one-loop bispectrum Bakx et al. (2025) and PNG inference from DESI data Chudaykin et al. (2025). We conclude that conservative scale cuts are essential to achieve unbiased inference of cosmological parameters.
@@ 5 Conclusion
Acknowledgments. MI thanks George Efstathiou for inspiring discussions. We thank Guido D’Amico for useful discussions. We are additionally grateful to the anonymous referee for useful comments, which helped improve the presentation. AC acknowledges funding from the Swiss National Science Foundation. All numerical calculations have been performed with the Helios cluster at the Institute for Advanced Study, Princeton. This work was supported in part by MEXT/JSPS KAKENHI Grant Numbers JP20H05861, JP21H01081, JP22K03634, JP24H00215 and JP24H00221.
Appendix A Analysis with $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$
We perform the BOSS-like analyses in the WC and EC2 models using a more conservative scale cut $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$ from the original analyses D’Amico et al. (2019); Colas et al. (2020).
We carried out a multi-chunk analysis of the four simulation samples using the BOSS data covariance estimated from the Patchy mocks. Additionally, we perform the analysis with the Gaussian covariance corresponding to 100 times the actual volume of each BOSS-like data chunks. We adopt $k_{\rm max}^{\rm z_{1}}=0.2\,h\text{Mpc}^{-1}$ for the $z_{1}$ redshift bin, as in the main analysis, but use a more conservative scale cut $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$ in the $z_{3}$ redshift bin. Further details of the analysis is given in Sec. 4.2.
Our results in the WC and EC2 models are presented in Fig. 4.
Figure 4: Posterior distributions inferred from the four BOSS-like samples of the PT Challenge simulation using the BOSS data covariance estimated from the Patchy mocks ($V_{\rm BOSS}$) and the analytical Gaussian covariance corresponding to 100 times the BOSS volume ($100\times V_{\rm BOSS}$). For $100\times V_{\rm BOSS}$, only the mean values of the corresponding parameters are depicted by dashed lines. For the WC and EC2 models, the $k_{\rm max}^{\rm z_{1}}/k_{\rm max}^{\rm z_{3}}=0.2/0.23\,h\text{Mpc}^{-1}$ are adopted, while for the EC3 model $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ is used for both redshift bins.
We also show the posteriors for the EC3 model which uses $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ for both $z_{1}/z_{3}$ redshift bins. The marginalized parameter constraints are listed in Tab. 3.
Param. Model
WC
$k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$
EC2
$k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$
$V_{\rm BOSS}$ $100\times V_{\rm BOSS}$ $V_{\rm BOSS}$ $100\times V_{\rm BOSS}$
$\Delta\omega_{cdm}/\omega_{cdm}$ $029_{-0.048}^{+0.044} 0.0$$15_{-0.005}^{+0.005} 0.0$$04_{-0.061}^{+0.053} 0.0$$0048_{-0.0067}^{+0.0069}$
$\Delta h/h$ $001_{-0.019}^{+0.018} 0.0$$017_{-0.0014}^{+0.0014} 0.0$$04_{-0.020}^{+0.019} 0.0$$0011_{-0.0017}^{+0.0017}$
$\Delta\sigma_{8}/\sigma_{8}$ $030_{-0.059}^{+0.055} 0.0$$38_{-0.005}^{+0.005} 0.0$$69_{-0.060}^{+0.053} 0.0$$037_{-0.005}^{+0.005}$
Table 3: Marginalized 1d constraints on the cosmological parameters from the BOSS-like analyses of the four simulation samples for the WC/EC2 models and two choices of the covariance matrix. All parameters are shown as $(p-p_{\rm fid.})/p_{\rm fid.}$, where $p_{\rm fid.}$ represents the fiducial value used in the simulations.
We found that the parameter constraints obtained with the scale cuts from Refs. D’Amico et al. (2019); Colas et al. (2020) are fully consistent with those from the baseline analysis employing $k_{\rm max}^{\rm z_{3}}=0.25\,h\text{Mpc}^{-1}$ , cf. Tab. 2. We identify a $4\%$ theory systematic bias in $\sigma_{8}$, which is slightly smaller than the $5\%$ bias observed in the main analysis. Thus, adopting $k_{\rm max}^{\rm z_{3}}=0.23\,h\text{Mpc}^{-1}$ does not illuminate the bias in the $\sigma_{8}$ recovery. We therefore conclude that the scale cuts $k_{\rm max}^{\rm z_{3}}=0.23$ and $0.25\,h\text{Mpc}^{-1}$ are overoptimistic and result in a significant systematic error in the estimation of $\sigma_{8}$. As shown in Sec. 4.1, this bias can be reduced by half when using $k_{\rm max}=0.2\,h\text{Mpc}^{-1}$ for both $z_{1}/z_{3}$ redshift bins in the EC2 model.